# Watching, Reasoning and Searching: A Video Deep Research Benchmark on Open Web for Agentic Video Reasoning

### v1

@@ 4.2 Main Results
From the overall averages, our VideoDR setting yields a clear capability stratification among existing models: Gemini-3-pro-preview leads under both paradigms (Workflow 69% / Agentic 76%), closely followed by GPT-5.2 which demonstrates comparable top-tier performance (Workflow 69% / Agentic 69%), while GPT-4o consistently forms the second tier (42% / 43%), and the three open-source models are overall weaker (Qwen3-Omni-30B-A3B 37% / 37%, InternVL3.5-14B 27% / 30%, MiniCPM-V 4.5 25% / 16%).
Difficulty stratification
reveals the essential differences between the two paradigms. As shown in Table 1, all evaluated models exhibit a highly consistent decline in performance as the difficulty level increases from Low to Mid and then to High. Human performance decreases from 90.00% to 50.56% and then to 10.63%, and models also generally degrade as difficulty increases, suggesting that our difficulty definition based on human success rate indeed corresponds to longer and more fragile evidence chains. Gains from Agentic mainly appear in models with stronger state maintenance capabilities, and become more pronounced toward Mid/High; conversely, mid-tier models can suffer a clear backlash on High. For example, Gemini-3-pro-preview continues to improve on Mid/High from 61.11%69.44% and 56.25%65.62%, and GPT-5.2 similarly achieves gains on High difficulty (62.50%→65.62%); but while GPT-4o improves on Low/Mid (50.00%62.50%, 30.56%38.89%), it drops sharply on High (46.88%28.12%). This phenomenon is highly consistent with our implementation constraints: in Agentic, the model cannot re-watch the video during research, so the subsequent process must rely solely on cues obtained from the initial viewing to guide multi-round search and reasoning, lacking opportunities for online correction via revisiting the video. Once early visual anchors drift, web noise can amplify the deviation in later searches, making High-difficulty questions more prone to goal drift; in contrast, Workflow externalizes video cues into an intermediate textual representation, effectively providing a repeatedly accessible external memory that reduces drift risk in long-horizon decision-making.
Video-duration stratification
further reinforces this explanation: as videos get longer, cues become more dispersed, and the state space grows, so Agentic places higher demands on a model’s ability to retain and continuously leverage initial video cues over the long run, leading to strong polarization. As shown in Table 2, Gemini-3-pro-preview achieves substantial gains on Medium/Long (65.79%84.21%, 50.00%70.00%), indicating that strong models can preserve enough discriminative details from one viewing and keep using them in subsequent search; in contrast, Qwen3-Omni-30B-A3B and MiniCPM-V 4.5 drop markedly on Long (50.00%20.00%, 30.00%10.00%), suggesting that when a model struggles to stably preserve and continuously utilize key video anchors across multiple rounds of search and reasoning, directly using the video does not translate into better search decisions and instead makes the system more likely to destabilize over long chains. In other words, long videos amplify the key trade-off between the two paradigms: Workflow may lose details but is more controllable, whereas Agentic preserves details more faithfully but relies more on long-term consistency.
Domain stratification
highlights the respective advantages of the two paradigms across domains. As shown in Table 3, in Technology, Gemini-3-pro-preview shows the largest Agentic-over-Workflow improvement (64.29%85.71%), indicating that such questions often require translating fine-grained visual cues into high-precision queries, and Workflow’s one-shot summary is more likely to miss crucial search triggers. However, in Geography, both Gemini-3-pro-preview (70.00%50.00) and GPT-4o (40.00%20.00) degrade noticeably, suggesting that geographic search is more ambiguous and therefore depends more on clear and consistent visual anchors to keep the search target stable.

### v2

@@ 4.2 Main Results
| Model | Workflow | Agentic | ||||||
| Low | Mid | High | Ave. | Low | Mid | High | Ave. | |
| #Samples | 112 | 171 | 217 | 500 | 112 | 171 | 217 | 500 |
| Qwen3.5-35B-A3B | 61.61 | 37.43 | 19.35 | 35.00 | 57.14 | 40.94 | 17.97 | 34.60 |
| InternVL3.5-14B | 29.46 | 16.96 | 8.29 | 16.00 | 30.36 | 16.96 | 6.45 | 15.40 |
| Gemma-4-31B-it | 52.68 | 38.60 | 15.21 | 31.60 | 57.14 | 43.27 | 13.82 | 33.60 |
| Gemini-3.1-pro-preview | 78.57 | 63.74 | 28.57 | 51.80 | 84.82 | 62.57 | 30.41 | 53.60 |
| GPT-5.4 | 66.96 | 52.63 | 32.26 | 47.00 | 71.43 | 53.80 | 29.95 | 47.40 |
| GLM-4.6V-Flash | 33.93 | 19.30 | 8.29 | 17.80 | 43.75 | 23.98 | 8.76 | 21.80 |
| Human | 81.71 | 49.64 | 11.39 | 40.22 | 81.71 | 49.64 | 11.39 | 40.22 |
| Model | Workflow | Agentic | ||||||
| Short | Medium | Long | Ave. | Short | Medium | Long | Ave. | |
| #Samples | 162 | 151 | 187 | 500 | 162 | 151 | 187 | 500 |
| Qwen3.5-35B-A3B | 30.86 | 35.76 | 37.97 | 35.00 | 31.48 | 33.77 | 37.97 | 34.60 |
| InternVL3.5-14B | 17.28 | 11.92 | 18.18 | 16.00 | 12.35 | 15.89 | 17.65 | 15.40 |
| Gemma-4-31B-it | 23.46 | 25.17 | 43.85 | 31.60 | 30.25 | 29.80 | 39.57 | 33.60 |
| Gemini-3.1-pro-preview | 51.85 | 49.01 | 54.01 | 51.80 | 54.32 | 47.02 | 58.29 | 53.60 |
| GPT-5.4 | 42.59 | 43.71 | 53.48 | 47.00 | 43.83 | 40.40 | 56.15 | 47.40 |
| GLM-4.6V-Flash | 15.43 | 18.54 | 19.25 | 17.80 | 20.37 | 23.84 | 21.39 | 21.80 |
| Human | 41.65 | 37.91 | 40.85 | 40.22 | 41.65 | 37.91 | 40.85 | 40.22 |
| Model | Setting | Domain (%) | ||||||
| History | Geography | Culture | Economics | Technology | Daily Life | Ave. | ||
| #Samples | 59 | 38 | 65 | 57 | 113 | 168 | 500 | |
| Qwen3.5-35B-A3B | Workflow | 32.20 | 28.95 | 32.31 | 22.81 | 30.09 | 45.83 | 35.00 |
| Agentic | 28.81 | 28.95 | 27.69 | 35.09 | 34.51 | 40.48 | 34.60 | |
| InternVL3.5-14B | Workflow | 18.64 | 5.26 | 15.38 | 12.28 | 15.93 | 19.05 | 16.00 |
| Agentic | 18.64 | 2.63 | 20.00 | 24.56 | 6.19 | 18.45 | 15.40 | |
| Gemma-4-31B-it | Workflow | 23.73 | 28.95 | 26.15 | 19.30 | 30.09 | 42.26 | 31.60 |
| Agentic | 27.12 | 18.42 | 32.31 | 24.56 | 37.17 | 40.48 | 33.60 | |
| Gemini-3.1-pro-preview | Workflow | 52.54 | 47.37 | 49.23 | 42.11 | 50.44 | 57.74 | 51.80 |
| Agentic | 59.32 | 47.37 | 49.23 | 35.09 | 54.87 | 60.12 | 53.60 | |
| GPT-5.4 | Workflow | 52.54 | 44.74 | 52.31 | 33.33 | 39.82 | 52.98 | 47.00 |
| Agentic | 47.46 | 55.26 | 46.15 | 29.82 | 46.02 | 52.98 | 47.40 | |
| GLM-4.6V-Flash | Workflow | 16.95 | 10.53 | 20.00 | 12.28 | 21.24 | 18.45 | 17.80 |
| Agentic | 20.34 | 10.53 | 27.69 | 21.05 | 18.58 | 25.00 | 21.80 | |
| Human | 39.62 | 37.97 | 47.81 | 39.53 | 37.94 | 39.77 | 40.22 | |
From the overall averages, our VideoDR setting yields a clear capability stratification among existing models. Gemini-3.1-pro-preview leads under both paradigms, achieving 51.80% under Workflow and 53.60% under Agentic. GPT-5.4 demonstrates the second-best overall performance, with 47.00% under Workflow and 47.40% under Agentic. Qwen3.5-35B-A3B and Gemma-4-31B-it form a middle tier, with Qwen3.5-35B-A3B achieving 35.00% / 34.60% and Gemma-4-31B-it achieving 31.60% / 33.60% under Workflow / Agentic, respectively. InternVL3.5-14B and GLM-4.6V-Flash are overall weaker, with InternVL3.5-14B reaching 16.00% / 15.40% and GLM-4.6V-Flash reaching 17.80% / 21.80%. Compared with human performance (40.22%), only Gemini-3.1-pro-preview and GPT-5.4 consistently surpass the human average under this evaluation setting.
Difficulty stratification. This reveals the essential differences between the two paradigms. As shown in Table 1, all evaluated models exhibit a highly consistent decline in performance as the difficulty level increases from Low to Mid and then to High. Human performance decreases from 81.71% to 49.64% and then to 11.39%, and models also generally degrade as difficulty increases, suggesting that our difficulty definition based on human success rate indeed corresponds to longer and more fragile evidence chains. However, the effect of Agentic varies substantially across models and difficulty levels. For example, Gemini-3.1-pro-preview improves on Low and High difficulty levels, from 78.57%84.82% and 28.57%30.41%, respectively, while slightly dropping on Mid from 63.74% to 62.57%. GPT-5.4 improves on Low/Mid (66.96%71.43%, 52.63%53.80%), but drops on High (32.26%29.95%). Qwen3.5-35B-A3B and InternVL3.5-14B also show declines on High when switching to Agentic, from 19.35%17.97% and 8.29%6.45%, respectively. In contrast, GLM-4.6V-Flash improves slightly on High from 8.29% to 8.76%, although its absolute performance remains low. These results indicate that directly placing the video inside the agent loop does not automatically improve difficult cases. This phenomenon can be partly explained by our implementation constraints: in Agentic, the model cannot re-watch the video during research, so the subsequent process must rely solely on cues obtained from the initial viewing to guide multi-round search and reasoning, lacking opportunities for online correction via revisiting the video. Once early visual anchors drift, web noise can amplify the deviation in later searches, making High-difficulty questions more prone to goal drift; in contrast, Workflow externalizes video cues into an intermediate textual representation, effectively providing a repeatedly accessible external memory that reduces drift risk in long-horizon decision-making.
Video-duration stratification. This further reinforces the above observation: as videos get longer, cues become more dispersed and the state space grows, so Agentic places higher demands on a model’s ability to retain and continuously leverage initial video cues, leading to strong polarization. As shown in Table 2, Gemini-3.1-pro-preview gains on Short and Long videos, improving from 51.85%54.32% and 54.01%58.29%, respectively, while dropping on Medium videos from 49.01% to 47.02%. GPT-5.4 shows a similar mixed pattern, improving on Short and Long videos (42.59%43.83%, 53.48%56.15%) but dropping on Medium videos (43.71%40.40%). For weaker or less stable models, the effect is also heterogeneous: Qwen3.5-35B-A3B slightly improves on Short videos, drops on Medium videos, and remains unchanged on Long videos; InternVL3.5-14B drops on Short and Long videos but improves on Medium videos; Gemma-4-31B-it improves on Short and Medium videos but drops on Long videos from 43.85% to 39.57%. GLM-4.6V-Flash improves across all duration groups, but its absolute accuracy remains much lower than that of the leading models. These results suggest that Agentic does not uniformly benefit from longer or richer video input: when a model struggles to preserve and reuse key video anchors across multiple rounds of search and reasoning, direct video access may fail to improve search decisions and can even destabilize long reasoning chains. In other words, long videos amplify the trade-off between the two paradigms: Workflow may lose some visual details but provides more controllable textual anchors, whereas Agentic has direct access to raw video information but depends more on long-term state consistency.
Domain stratification. This highlights the respective advantages of the two paradigms across domains. As shown in Table 3, Gemini-3.1-pro-preview benefits from Agentic in History, Technology, and Daily Life, improving from 52.54%59.32%, 50.44%54.87%, and 57.74%60.12%, respectively. It remains unchanged in Geography and Culture, while dropping in Economics from 42.11% to 35.09%. GPT-5.4 shows a different pattern: it improves in Geography and Technology (44.74%55.26%, 39.82%46.02%), remains unchanged in Daily Life, but drops in History, Culture, and Economics. Other models also exhibit domain-specific trade-offs. For instance, Gemma-4-31B-it improves in History, Culture, Economics, and Technology, but drops in Geography and Daily Life; Qwen3.5-35B-A3B improves in Economics and Technology, but drops in History, Culture, and Daily Life. This domain-level variation further shows that the benefit of Agentic is not universal, but depends jointly on the model’s video-memory stability and the domain-specific ambiguity of downstream search. Domains that require fine-grained visual cues to be translated into precise search constraints may benefit from direct video access when the underlying model is strong enough, whereas domains with ambiguous or weak visual anchors remain vulnerable to search drift.
| Model | Setting | Error Type Count | ||||||||
| Qwen3.5-35B-A3B | Workflow | 129 | 22 | 21 | 110 | 20 | 10 | 13 | 0 | 325 |
| Agentic | 114 | 25 | 24 | 114 | 35 | 10 | 5 | 0 | 327 | |
| InternVL3.5-14B | Workflow | 155 | 31 | 56 | 141 | 15 | 14 | 6 | 2 | 420 |
| Agentic | 150 | 26 | 32 | 156 | 26 | 17 | 7 | 9 | 423 | |
| Gemma-4-31B-it | Workflow | 115 | 1 | 95 | 91 | 16 | 10 | 2 | 12 | 342 |
| Agentic | 118 | 6 | 76 | 95 | 4 | 15 | 10 | 8 | 332 | |
| Gemini-3.1-pro-preview | Workflow | 79 | 27 | 15 | 84 | 13 | 17 | 4 | 2 | 241 |
| Agentic | 86 | 16 | 4 | 95 | 15 | 13 | 3 | 0 | 232 | |
| GPT-5.4 | Workflow | 103 | 17 | 15 | 109 | 9 | 8 | 2 | 2 | 265 |
| Agentic | 99 | 17 | 11 | 123 | 4 | 8 | 1 | 0 | 263 | |
| GLM-4.6V-Flash | Workflow | 15 | 52 | 53 | 106 | 42 | 5 | 101 | 37 | 411 |
| Agentic | 25 | 58 | 38 | 109 | 41 | 7 | 96 | 17 | 391 | |
4.3 Error Analysis
Perception Grounding Limitation. Table 4 indicates that Categorical Error remains a major failure mode for many models, especially Qwen3.5-35B-A3B, InternVL3.5-14B, and Gemma-4-31B-it, while Numerical Error is comparably prominent for stronger models such as Gemini-3.1-pro-preview and GPT-5.4. Under Workflow, categorical errors range from 79 to 155 among these representative models, and under Agentic, they remain high, ranging from 86 to 150. This pattern is closely tied to our setting: once the research phase cannot revisit the video, downstream retrieval and reasoning depend heavily on the visual anchors extracted from the first pass. When early perception deviates, there is no mechanism to re-localize key frames and correct the anchor, making error propagation along the evidence chain more likely. At the same time, the error distribution also reveals model-specific weaknesses. For example, GLM-4.6V-Flash has an unusually large number of Reasoning Errors under both Workflow and Agentic settings (101 and 96), indicating that its failures are not limited to perception grounding, but also involve evidence integration and reasoning instability. This further suggests that different models fail at different stages of the overall video-grounded open-domain QA pipeline.
Numerical Error Bottleneck. Table 4 shows that Gemini-3.1-pro-preview has the lowest overall error count among all models, with 241 total errors under Workflow and 232 under Agentic, yet this advantage does not eliminate Numerical Error. Under Workflow, Gemini-3.1-pro-preview still has 84 numerical errors, while Qwen3.5-35B-A3B, InternVL3.5-14B, Gemma-4-31B-it, GPT-5.4, and GLM-4.6V-Flash have 110, 141, 91, 109, and 106, respectively. Under Agentic, Gemini-3.1-pro-preview records 95 numerical errors, compared with 114, 156, 95, 123, and 109 for the same five models. Numerical errors therefore remain substantial across all evaluated models, even for the strongest systems. This suggests that numerical reliability constitutes a distinct and persistent weakness for current MLLMs: while stronger models can reduce not-found errors and improve evidence grounding, they still frequently fail when the task requires precise counting, comparison, or numerical verification.
5 Conclusion
Appendix A Retrieval-and-Reasoning Analysis
As shown in Table 5, the number of search calls and reasoning steps is not monotonically correlated with accuracy. What determines performance is not using search more, but whether a model can turn a limited number of search calls into a high-quality evidence chain. Concretely, Gemini-3.1-pro-preview demonstrates stronger retrieval-and-reasoning effectiveness: under the Agentic setting, it averages 4.55/3.16 reasoning/search calls with 345.73s runtime and achieves the best overall score of 53.60%, indicating that additional retrieval and reflection are converted into more reliable evidence integration for this model. In contrast, GLM-4.6V-Flash consumes the longest runtime under both paradigms, with 511.98s in Workflow and 502.96s in Agentic, but only achieves 17.80% / 21.80% accuracy, suggesting that longer execution time does not necessarily lead to better evidence acquisition or answer correctness. Qwen3.5-35B-A3B also shows that reducing reasoning and search calls does not automatically improve or harm performance in a simple way: its Agentic setting uses fewer reasoning/search calls than Workflow, 1.59/3.02 vs. 2.97/3.80, and runs faster, 374.47s vs. 485.57s, but accuracy still slightly drops from 35.00% to 34.60%. Gemma-4-31B-it provides another contrasting case: Agentic improves accuracy from 31.60% to 33.60%, while increasing reasoning calls from 3.69 to 3.84, reducing search calls from 2.00 to 1.70, and increasing runtime from 232.28s to 285.67s. These observations suggest that interaction quality, rather than raw interaction frequency or runtime, is the key factor determining whether additional search and reflection can improve performance.
| Model | Workflow | Agentic | ||||
| reasoning | search | time (s) | reasoning | search | time (s) | |
| Qwen3.5-35B-A3B | 2.97 | 3.80 | 485.57 | 1.59 | 3.02 | 374.47 |
| InternVL3.5-14B | 2.22 | 2.84 | 154.03 | 1.86 | 2.49 | 126.37 |
| Gemma-4-31B-it | 3.69 | 2.00 | 232.28 | 3.84 | 1.70 | 285.67 |
| Gemini-3.1-pro-preview | 3.58 | 2.30 | 188.20 | 4.55 | 3.16 | 345.73 |
| GPT-5.4 | 3.88 | 2.86 | 178.84 | 3.84 | 2.78 | 166.32 |
| GLM-4.6V-Flash | 2.05 | 2.40 | 511.98 | 1.92 | 2.39 | 502.96 |
Appendix B Token Usage Analysis
As shown in Table 6, Agentic generally introduces higher token consumption than Workflow, especially in prompt tokens, because the model directly processes the raw video-conditioned context and autonomously maintains the retrieval-and-reasoning trajectory within a single execution loop. This trend is particularly evident for Qwen3.5-35B-A3B, whose average total tokens increase from 168,483 to 301,187, and for Gemini-3.1-pro-preview, whose average total tokens increase from 80,271 to 199,008. GLM-4.6V-Flash also shows a substantial increase from 56,839 to 96,329 total tokens. In contrast, GPT-5.4 is the only model whose Agentic setting slightly reduces average total token usage, from 109,236 to 101,936, suggesting more compact interaction behavior under the end-to-end agent loop. Across models, completion tokens remain relatively stable or even decrease in several Agentic settings, while prompt tokens dominate the overall increase, indicating that the main efficiency burden of Agentic comes from maintaining larger contextual states rather than generating longer responses. The high P95 values, especially for Qwen3.5-35B-A3B under Agentic, further suggest that Agentic can occasionally trigger very long search-and-reasoning trajectories, making token cost less predictable than Workflow.
| Model | Setting |
| Qwen3.5-35B-A3B | Workflow | 159621 | 8861 | 168483 | 163890 | 337718 | ||||||||||
| Agentic | 295004 | 6183 | 301187 | 200927 | 755463 | |||||||||||
| InternVL3.5-14B | Workflow | 53209 | 7481 | 60690 | 53202 | 123752 | ||||||||||
| Agentic | 59108 | 6379 | 65487 | 58544 | 131373 | |||||||||||
| Gemma-4-31B-it | Workflow | 54814 | 5458 | 60272 | 37538 | 174100 | ||||||||||
| Agentic | 70917 | 5676 | 76592 | 60790 | 195298 | |||||||||||
| Gemini-3.1-pro-preview | Workflow | 71949 | 8322 | 80271 | 66936 | 153754 | ||||||||||
| Agentic | 189733 | 9275 | 199008 | 153038 | 364158 | |||||||||||
| GPT-5.4 | Workflow | 103461 | 5775 | 109236 | 99682 | 187596 | ||||||||||
| Agentic | 96933 | 5003 | 101936 | 94815 | 169805 | |||||||||||
| GLM-4.6V-Flash | Workflow | 50289 | 6550 | 56839 | 47409 | 132597 | ||||||||||
| Agentic | 90545 | 5784 | 96329 | 83331 | 203463 |
Appendix C Efficiency Analysis
As shown in Table 7, higher accuracy does not necessarily imply higher efficiency, and the relative efficiency of Workflow and Agentic varies substantially across models. Gemini-3.1-pro-preview achieves the best accuracy under Agentic, improving from 51.80% to 53.60%, but this gain comes with much higher average runtime, increasing from 188.20s to 345.73s, and lower token efficiency, dropping from 0.65 to 0.27 accuracy per 1K tokens. A similar efficiency trade-off appears for Gemma-4-31B-it, where Agentic improves accuracy from 31.60% to 33.60% but increases both runtime and tokens per correct sample. By contrast, GPT-5.4 shows the most favorable Agentic efficiency profile and remains relatively stable across the two paradigms: Agentic slightly improves accuracy from 47.00% to 47.40%, while reducing average runtime from 178.84s to 166.32s and improving both accuracy per 1K tokens and accuracy per minute. For weaker models, Agentic often improves speed but not necessarily cost-effectiveness; for example, Qwen3.5-35B-A3B runs faster under Agentic but requires far more tokens per correct sample, increasing from 481,379 to 870,483. Overall, these results reinforce that Agentic is not uniformly more efficient than Workflow: its practical value depends on whether the additional context and autonomous reasoning process can be converted into accuracy gains without excessive token or latency overhead.
| Model | Setting |
| Qwen3.5-35B-A3B | Workflow | 35.00 | 485.57 | 382.97 | 1300.26 | 0.21 | 4.32 | 481378.61 | ||||||||||||||
| Agentic | 34.60 | 374.47 | 284.48 | 991.93 | 0.11 | 5.54 | 870482.53 | |||||||||||||||
| InternVL3.5-14B | Workflow | 16.00 | 154.03 | 124.35 | 349.50 | 0.26 | 6.23 | 379312.29 | ||||||||||||||
| Agentic | 15.40 | 126.37 | 111.28 | 303.93 | 0.24 | 7.31 | 425239.98 | |||||||||||||||
| Gemma-4-31B-it | Workflow | 31.60 | 232.28 | 172.94 | 567.54 | 0.52 | 8.16 | 190733.88 | ||||||||||||||
| Agentic | 33.60 | 285.67 | 197.42 | 854.56 | 0.44 | 7.06 | 227953.85 | |||||||||||||||
| Gemini-3.1-pro-preview | Workflow | 51.80 | 188.20 | 149.44 | 365.58 | 0.65 | 16.51 | 154963.72 | ||||||||||||||
| Agentic | 53.60 | 345.73 | 300.30 | 783.26 | 0.27 | 9.30 | 371283.09 | |||||||||||||||
| GPT-5.4 | Workflow | 47.00 | 178.84 | 139.70 | 356.98 | 0.43 | 15.77 | 231414.84 | ||||||||||||||
| Agentic | 47.40 | 166.32 | 131.35 | 339.31 | 0.47 | 17.10 | 214491.07 | |||||||||||||||
| GLM-4.6V-Flash | Workflow | 17.80 | 511.98 | 402.86 | 1224.30 | 0.31 | 2.09 | 319319.05 | ||||||||||||||
| Agentic | 21.80 | 502.96 | 362.78 | 1283.71 | 0.23 | 2.60 | 441875.81 |
Appendix D Annotation Effort
Constructing VideoDR requires substantial manual effort because each sample must satisfy both video dependency and web dependency. Annotators need to inspect videos repeatedly, identify multi-frame visual anchors, design questions that require multi-hop open-web verification, archive supporting evidence, and further filter out samples answerable by video-only or web-only shortcuts. Although this process is time-consuming and limits rapid scaling, it is necessary to ensure that each retained question has a unique, verifiable answer and genuinely evaluates video-grounded deep research rather than simple video understanding or text-based web search.
Appendix E Data
We also prepare an open-source dataset release at:
The dataset release contains the benchmark annotations and the structured fields needed to reconstruct benchmark samples, including the question-answer pairs, video references, evidence references, and sample-level metadata used for evaluation.
To facilitate reproducibility and long-term auditing, we preserve local copies of all videos and evidence webpages involved in annotation and verification. Each benchmark item is associated with source references and archived evidence records. In the public release, we focus on benchmark annotations, metadata, source URLs, evidence references, and evaluation scripts. Third-party raw videos or webpages are redistributed only when their licenses or terms of use explicitly permit such redistribution.
Appendix F Previous Version Experimental Results
This section reports the experimental tables from the previous version (v1) of this manuscript.
F.1 Overall Performance by Difficulty
Table 8 reports the v1 accuracy results under the Workflow and Agentic settings across Low, Mid, and High difficulty groups. In the v1 setting, difficulty was also defined according to human success rate, and the table was used to analyze how model performance changes as the evidence chain becomes longer and more fragile.
@@ F.1 Overall Performance by Difficulty
F.2 Performance by Video Duration
Table 9 reports the v1 results grouped by video duration. This table studyed whether longer videos make it harder for models to preserve and reuse the initial visual anchors during subsequent web search and reasoning.
@@ F.2 Performance by Video Duration
F.3 Performance by Domain
Table 10 reports the v1 domain-wise results. The table compares Workflow and Agentic performance across History, Geography, Culture, Economy, Technology, and Daily Life questions, and was used to analyze whether different semantic domains lead to different search-drift patterns.
@@ F.3 Performance by Domain
F.4 Tool-use Statistics
Table 11 reports the v1 retrieval-and-reasoning statistics, including average think calls, search calls, and runtime under both paradigms. These statistics were used to examine whether stronger performance came from more frequent tool use or from more effective evidence acquisition and filtering.